# Semantic-guided Search for Efficient Program Repair with Large Language Models

### v1

@@ 4.5. RQ5: Effectiveness of semantic-guided patch generation
LLM Beam Search Multiple Sampling FLAMES
CodeGen-2B 38 71 123
InCoder-1B 96 88 138
RepairLLama 75 137 210
CodeGen-6B 11 71 137
InCoder-6B 27 87 144
In this experiment, we assess the effectiveness of FLAMES’s semantic-guided patch generation versus beam search and multiple sampling. We evaluated these methods with various LLMs: CodeGen-2B, InCoder-1B, RepairLlama, CodeGen-6B, and InCoder-6B, measuring effectiveness by the number of plausible patches. Detailed results are in Table 5.
FLAMES consistently outperforms beam search across all models, with improvement rates ranging from 43% to 1145%. For instance, with RepairLlama, FLAMES generates 210 plausible patches, nearly three times the 75 produced by beam search. This is due to FLAMES ’s effective use of GPU resources which allows for a more thorough exploration of the large patch space. In contrast, beam search requires substantial VRAM, resulting in a high out-of-memory crash rate as discussed in RQ4 (Section 4.4).
We also observes that multiple sampling can significantly improve upon beam search across various models. This approach involves using beam search with a low beam size and then sampling the model multiple times to generate additional candidates. As a result, it can partially mitigate the memory inefficiencies of beam search and scale the patch space. Despite this, FLAMES still outperforms multiple sampling across models by 53% to 93%. We suspect that this is because our patch generation process is guided by feedback from test validation, whereas multiple sampling are purely random based on model’s outputs.
In conclusion, the consistent performance improvements across various models strongly support the effectiveness of FLAMES’s semantic-guided patch generation compared to traditional beam search and multiple sampling.
Answers to RQ5 (Effectiveness of Patch Generation): FLAMES’s semantic-guided patch generation demonstrates superior performance across multiple LLMs in APR tasks, consistently outperforming widely-used patch generation methods: beam search and multiple sampling by at least 43%.
@@ 4.6. RQ6: Ablation Study
To systematically assess how each component influence the performance of FLAMES, we investigated three key factors: the semantic-guided reward function (for evaluation phase), the size of the expansion $k$ (for expansion phase), and the choice of policy algorithm (for selection phase).
@@
In this experiment, we evaluate the effectiveness of our reward function, as detailed in Section 3.5. We hypothesize that this function will guide the generation of more plausible patches by leveraging the ratio of passing test cases. To test this hypothesis, we compare our approach with a variant that depends solely on the models’ probability outputs and disregards feedback from our semantic reward function. Our experimental results showed that incorporating a semantic reward function can lead to a 40% increase, from 150 to 210, of plausible patches . This substantial improvement highlights the value of semantic guidance in refining the search process to concentrate on patches that is more likely to succeed.
@@
Next, we explore how the expansion size $k$ affects FLAMES. To investigate this, we conducted experiments with various values of $k$, including 3, 5, 7, and 10. Our findings show that as $k$ increased from 3 to 10, the number of plausible patches rose from 190 to 210. Correspondingly, the average time required per patch also increased, from 49.19 seconds when $k=3$ to 98.41 seconds when $k=10$. This trend suggests that a larger $k$ allows for a more thorough exploration of the solution space, which enhances the likelihood of identifying high-quality patches. However, this comes at the expense of increased computational time, highlighting a trade-off between depth of exploration and performance efficiency.
@@ 5.1. Limitations
First, while our approach, FLAMES, correctly fixes 21 more bugs and 42 unique bugs compared to the original RepairLlama, there are 21 bugs that the original model could repair but FLAMES could not. In a deeper analysis, we found that this issue is primarily due to the relatively small expansion size ($k=10$) used in our experiments. This limitation compromise FLAMES’s exploration, resulting in a less comprehensive search space than that of RepairLlama’s beam search, which considers up to 25 possible tokens of expansions. Consequently, FLAMES misses some bugs that could potentially be fixed with a broader search scope. Although a straightforward solution would be to increase the expansion size $k$, doing so introduces a trade-off with time efficiency, as shown in Section 4.6.2. Therefore, further analysis is essential to optimize the balance between expansion size and efficiency to enhance FLAMES’s performance.
@@ 5.2. Threats to validity
A main threat to internal validity stems from the issue of data leakage, where evaluation data might overlap with training and fine-tuning data. Following prior studies (Silva et al., 2023; Jiang et al., 2023a), we employed the HumanEval-Java dataset for evaluation, which comprises 163 artificially introduced bugs within the HumanEval benchmark (Chen et al., 2021c), thereby minimizing the risk of data leakage. Another concern is the manual assessment of patch correctness. To address this, authors and an external annotator rigorously analyze each patch to confirm semantic equivalence to developer-written patches. Moreover, to ensure transparency, we also have made all patches generated by NPR techniques and our assessment results available in our replication package.
@@
Our research shares a common threat to external validity with prior works (Jiang et al., 2021; Silva et al., 2023; Xia and Zhang, 2022; Xia et al., 2023d; Wei et al., 2023; Le-Cong et al., 2024) that our results may not generalize across different programming languages and datasets. To mitigate this risk, we evaluated our model, FLAMES, and state-of-the-art baselines using two well-known datasets, Defects4J (Just et al., 2014) and HumanEval-Java (Jiang et al., 2023a), and found our performance robust across these datasets. Additionally, the core contributions of our approach are language-agnostic, suggesting potential applicability to various programming languages. Therefore, we believe this risk to be minimal.
@@
Our methodology may encounter threats to construct validity concerning the appropriateness of our evaluation metrics. Due to limited human resources, beside the number of correct patches, we leverages the number of plausible patches as as a proxy effectiveness metric, which might not fully capture the actual effectiveness of APR techniques. To minimize this risk, we only utilized this metric when comparing different variants of the same APR techniques, where we observed consistent trends between correct patches and plausible patches.
@@ 6.1. Large Language Models for Automated Program Repair
Recent advancements in Large Language Models (LLMs), pre-trained on extensive datasets, have significantly advanced various coding tasks (Fan et al., 2023), such as code generation (Nijkamp et al., 2022; Fried et al., 2022; Chen et al., 2021c), understanding (Nam et al., 2024; Ahmed and Devanbu, 2022; Zhou et al., 2024), and analysis (Li et al., 2024; Kazerounian et al., 2021; Le-Cong et al., 2022). In the field of APR, LLMs have also shown great promise when employed through either Fine-tuning techniques (Silva et al., 2023; Jiang et al., 2023a; Huang et al., 2023; Paul et al., 2023; Zhang et al., 2022; Zirak and Hemmati, 2024) or Prompting methods (Ahmed and Devanbu, 2023; Xia et al., 2023d; Xia and Zhang, 2023).
@@
Different from these studies that typically employ beam search to generate candidate patches, our work empirically investigates the limitations of beam search, particularly when implemented on standard hardware setups. Motivated by these shortcomings, we propose the use of semantic-guided best-first search to guide the repair process of LLMs.
@@ 6.2. Efficiency of Automated Program Repair
In addition to evaluating effectiveness, several studies also focus on the efficiency of APR (Liu et al., 2020; Chen et al., 2021b; Qi et al., 2013; Xiao et al., 2024; Chen et al., 2022). Liu et al. (Liu et al., 2020) conducted a systematic assessment of 16 APR techniques, revealing inefficiencies in state-of-the-art APR tools and advocating for further industry exploration of efficiency impacts. Chen et al. (Chen et al., 2021b) introduced UniAPR, a patch validation framework that enhances efficiency by avoiding unnecessary restarts of the JVM for all existing bytecode and source code-level APR techniques. Xiao et al. (Xiao et al., 2023) developed ExpressAPR, which accelerates APR by integrating five mechanisms: mutant schemata, mutant deduplication, test virtualization, test prioritization, and parallelization. Chen et al. (Chen et al., 2022) proposed utilizing XGBoost to develop on-the-fly prioritization techniques to expedite APR techniques. Kim et al. proposed CASINO (Kim et al., 2023) and Gresino (Kim et al., 2024), novel patch-scheduling algorithms for enhancing time efficiency of APR techniques.
@@ 7. Conclusion
In this study, we empirically evaluated the impact of beam size on the memory efficiency and effectiveness of LLM-based APR techniques. Our findings reveal that an increase in beam size results in significantly higher VRAM usage. Consequently, increasing the beam size cause numerous out-of-memory crashes and subsequently unduly degrades the performance of LLM-based APR. To address this challenge, we introduced FLAMES —the first APR approach that integrates LLMs with a semantic-guided best-first algorithm to guide the repair process. Our evaluation on the Defects4J and HumanEval-Java datasets demonstrates that FLAMES achieves state-of-the-art results. Additionally, FLAMES significantly reduces memory consumption by up to 83% and accelerates the repair process compared to conventional LLM-based APR techniques.
8. Data Availability.
We published our GitHub repository at https://github.com/thanhlecongg/FLAMES and replication package including dataset and results and appendix at  (Authors, 2023).

### v2

@@ 4.5. RQ5: Effectiveness of Semantic Guided Patch Generation
LLM BS MS FLAMES
CodeGen-2B 41 38 57
InCoder-1B 57 52 64
RepairLLama 112 90 133
CodeGen-6B 49 39 70
InCoder-6B 50 44 74
Qwen2.5-Coder-32B _ 68 91
A common approach to mitigate memory constraints is multiple sampling, which involves performing beam search with a low beam size and subsequently sampling the model multiple times to generate additional candidate patches. This method enables the generation of more candidate patches while reducing memory consumption. However, our findings indicate that this approach results in lower effectiveness compared to beam search. For instance, in the case of RepairLLama, multiple sampling leads to an approximately 20% decline in performance. Consequently, FLAMES significantly outperforms multiple sampling across models, achieving improvements ranging from 23% to 68%. We hypothesize that this performance gap arises because FLAMES leverages feedback from test validation to guide patch generation, whereas multiple sampling relies solely on random sampling from the model’s output.
In conclusion, consistent performance improvements across various models strongly support the effectiveness of FLAMES’s semantic-guided patch generation compared to widely used patch generation approaches such as beam search and multiple sampling.
Answers to RQ5 (Effectiveness of Patch Generation): FLAMES’s semantic-guided patch generation demonstrates superior performance across multiple LLMs in APR tasks, consistently outperforming widely used patch generation methods: beam search and multiple sampling, by at least 12% and 23%, respectively.
@@ 4.6. RQ6: Ablation study
To systematically assess how each component influences the performance of FLAMES, we investigated three key factors: the semantic-guided reward function (for the evaluation phase), the size of the expansion $k$ (for the expansion phase), and the choice of the policy algorithm (for the selection phase).
@@
In this experiment, we evaluate the effectiveness of our reward function, as detailed in Section 3.4. We hypothesize that this function will guide the generation of more plausible patches by leveraging the ratio of passing test cases. To test this hypothesis, we compare our approach with a variant that depends solely on the models’ probability outputs and disregards feedback from our semantic reward function. Our experimental results show that incorporating a semantic reward function can lead to a 40% increase, from 150 to 210, in plausible patches. This substantial improvement highlights the value of semantic guidance in refining the search process to concentrate on patches that are more likely to succeed.
@@
Next, we explore how the expansion size $k$ affects FLAMES by experimenting with various values of $k$, including 3, 5, 7, 10, 15, 20, 25 and 30. Our findings show that as $k$ increased from 3 to 10, the number of plausible patches increased from 190 to 210. Unfortunately, a further increase of $k$ to up to 30 showed minimal gains, with plausible patches ranging from 202 to 204. This is due to the broader, yet shallower search at higher k, which favors exploration over the exploitation needed for certain bugs. This trend suggests that a larger $k$ may allow for a more thorough exploration of the solution space, enhancing the likelihood of identifying high-quality patches. However, this comes at a trade-off between the depth of exploration and exploitation.
@@ 5.1. Limitations
First, while our approach, FLAMES, correctly fixes 21 more bugs and 42 unique bugs compared to the original RepairLlama, there are 21 bugs that the original model could repair but FLAMES could not. In an in-depth analysis on the 21 bugs that FLAMES cannot repair, we observed that the frequent pattern that these bugs most often require entirely new code segments. However, we also notice that these patterns appear in bugs that can be successfully repaired by FLAMES but cannot be repaired by RepairLlama. Therefore, it does not show a strong separation between the patterns of bugs fixed and not fixed by FLAMES. Therefore, we hypothesize that the failures stem from FLAMES’s trade-off strategy rather than specific types of bug. Specifically, FLAMES’s failure to repair these bugs can possibly be attributed to FLAMES’s smaller expansion size of 10. This limitation compromises FLAMES’s exploration, resulting in a less comprehensive search space than that of RepairLlama’s beam search, which considers up to 25 possible tokens of expansions. Consequently, FLAMES misses some bugs that could potentially be fixed with a broader search scope. Although a straightforward solution would be to increase the expansion size $k$, doing so introduces a trade-off between exploitation and exploration, as shown in Section 4.6.2. Therefore, further analysis is essential to optimize the balance between expansion size and efficiency to enhance FLAMES’s performance.
@@ 5.2. Threats to validity
A main threat to internal validity stems from the issue of data leakage, where evaluation data might overlap with training and fine-tuning data. To mitigate this risk, we followed prior studies (Silva et al., 2023; Jiang et al., 2023a) and utilized the HumanEval-Java dataset for evaluation. This dataset consists of 163 bugs introduced by injecting faults into the HumanEval benchmark (Chen et al., 2021c). Additionally, we employed TransformedD4J, which comprises 1,090 bugs generated through semantic-preserving transformations of Defects4J, further reducing the likelihood of data leakage. Note that, HumanEval-Java and TransformedD4J datasets were publicly released in July 2023 and January 2024, respectively. In contrast, CodeLlama, the base LLM underlying RepairLlama, was released in August 2024. This temporal gap significantly reduces the likelihood of data leakage during the pre-training phase. Furthermore, the fine-tuning data used in the development of RepairLlama was extracted from the Boa dataset (Dyer et al., 2013) collected in September 2015, further minimizing the possibility of overlap with our evaluation datasets. Therefore, we believe that the risk of data leakage in our study is negligible. Another concern is the manual evaluation of patch correctness. To address this, authors and an external annotator rigorously analyze each patch to confirm semantic equivalence to developer-written patches. Moreover, to ensure transparency, we also have made all patches generated by NPR techniques and our assessment results available in our replication package.
@@
Our research shares a common threat to external validity with prior works (Jiang et al., 2021; Silva et al., 2023; Xia and Zhang, 2022; Xia et al., 2023d; Wei et al., 2023; Le-Cong et al., 2025) that our results may not generalize across different programming languages and datasets. To mitigate this risk, we evaluated our model, FLAMES, and state-of-the-art baselines using three datasets, Defects4J (Just et al., 2014), HumanEval-Java (Jiang et al., 2023a) and TransformedD4J (Le-Cong et al., 2025), and found our performance generalize across these datasets. Additionally, the core contributions of our approach are language-agnostic, suggesting potential applicability to various programming languages. Therefore, we believe that this risk is minimal.
Another external validity relates to our generalization to larger LLMs such as ChatGPT or Claude models. FLAMES modify the LLM decoding strategy and thus require access to LLMs. This avoids evaluation on such closed-source LLMs. As an alternative, we have done our best to conduct experiments on CodeQwen2.5, the best open source code LLM across leaderboards, and demonstrate the effectiveness of FLAMES on this model in Table 7. Moreover, our experiments for RQ5 demonstrate the generalizability of FLAMES across six well-known LLMs with diverse model architectures and sizes. Therefore, we believe that FLAMES is capable of generalizing effectively to different LLMs.
@@
Our methodology may encounter threats to construct validity concerning the appropriateness of our evaluation metrics. Due to limited human resources, besides the number of correct patches, we leverage the number of plausible patches as a proxy effectiveness metric, which might not fully capture the actual effectiveness of APR techniques. To minimize this risk, we only utilized this metric when comparing different variants of the same APR techniques, where we observed consistent trends between correct patches and plausible patches.
In addition, we also share a common threat to construct validity with previous work (Jiang et al., 2023a, 2021; Chen et al., 2021a; Wei et al., 2023; Xia and Zhang, 2022) on the assumptions of perfect fault localization. However, we did our best efforts to ensure a fair comparison between FLAMES and LLM-based Program Repair by adopting the same fault localization assumptions as our base models as we reused their preprocessing pipelines. For example, in the default setting, our base model, RepairLlama, requires perfect fault localization, so that FLAMES also needs to use perfect fault localization. However, for multi-location bugs, RepairLlama only needs a coarse-grained bug region. As shown in Section 4.3.2, FLAMES can effectively repair multi-location bugs under this assumption. These results suggest that FLAMES may operate without perfect fault localization, provided that its base models support such flexibility.
Finally, as FLAMES relies on test case, it is possible that its performance may not generalize to lower-quality test suite. To address this concern, we evaluated the impact of test suite quality on FLAMES by grouping bugs according to line coverage of their test suites: ¡60%, 60–80%, and ¿80%. FLAMES produced plausible patches for 65%, 69%, and 62% of bugs in these groups, and correct patches for 40%, 42%, and 40%, respectively. However, the precision for the ¿80% coverage group was slightly higher at 0.64, compared to 0.61 and 0.62 in the lower coverage groups. These results suggest that FLAMES performs consistently across varying test suite quality, with a slight decrease in precision on lower coverage suites.
@@ 6.1. Large Language Models for Automated Program Repair
Recent advancements in Large Language Models (LLMs), pre-trained on extensive datasets, have significantly advanced various coding tasks (Fan et al., 2023), such as code generation (Fried et al., 2022; Chen et al., 2021c), understanding (Nam et al., 2024; Ahmed and Devanbu, 2022), and analysis (Kazerounian et al., 2021; Le-Cong et al., 2022). In the field of APR, LLMs have also shown great promise when employed through either fine-tuning techniques (Silva et al., 2023; Jiang et al., 2023a; Huang et al., 2023) or prompting methods (Xia et al., 2023d; Xia and Zhang, 2023).
@@
Different from these studies that typically employ beam search to generate candidate patches, our work empirically investigates the limitations of beam search, particularly when implemented on standard hardware setups. Motivated by these shortcomings, we propose the use of PG-TD to guide LLM-based APR.
@@ 6.2. Efficiency of Automated Program Repair
In addition to evaluating effectiveness, several studies also focus on the efficiency of APR (Liu et al., 2020; Chen et al., 2021b; Qi et al., 2013; Xiao et al., 2024; Chen et al., 2022). Liu et al. (Liu et al., 2020) conducted a systematic assessment of 16 APR techniques, revealing inefficiencies in state-of-the-art APR tools and advocating for further industry exploration of efficiency impacts. Chen et al. (Chen et al., 2021b) introduced UniAPR, a patch validation framework that enhances efficiency by avoiding unnecessary restarts of the JVM for all existing bytecode and source code-level APR techniques. Xiao et al. (Xiao et al., 2023) developed ExpressAPR, which accelerates APR by integrating five mechanisms: mutant schemata, mutant deduplication, test virtualization, test prioritization, and parallelization. Chen et al. (Chen et al., 2022) proposed utilizing XGBoost to develop on-the-fly prioritization techniques to expedite APR. Kim et al. (Kim et al., 2023) and Gresino (Kim et al., 2024) presented novel patch-scheduling algorithms for enhancing APR time efficiency.
@@ 7. Conclusion
In this study, we empirically evaluated the impact of beam size on the memory efficiency and effectiveness of LLM-based APR techniques. Our findings reveal that an increase in beam size results in significantly higher VRAM usage. Consequently, increasing the beam size causes numerous out-of-memory crashes and subsequently unduly degrades the performance of LLM-based APR. To address this challenge, we introduced FLAMES —the first APR approach that integrates LLMs with a semantic-guided best-first algorithm to guide the repair process. Our evaluation on various datasets demonstrates that FLAMES substantially outperforms the state-of-the-art baselines. Additionally, FLAMES significantly reduces memory consumption by up to 83% and accelerates the repair process compared to conventional LLM-based APR techniques.
Data Availability. We published replication package including dataset and results and appendix at (Authors, 2023).